# Imprints of quantum gravity effects on gravitational waves: a comparative study using extreme mass-ratio inspirals

### v1

@@ Abstract
Within a generally covariant Hamiltonian framework of loop quantum gravity (LQG), two black hole models parameterized by a quantum correction $\zeta$ have recently been constructed. Using extreme mass-ratio inspirals (EMRIs) as high-precision probes, we investigate the imprints of this LQG deformation in the surrounding spacetime. Waveforms generated via an improved augmented analytic kludge (AAK) model in both LQG-BH backgrounds and in Schwarzschild spacetime are compared through a faithfulness analysis. This allows us to quantify the detectability of the deviation with LISA and to derive constraints on $\zeta$ based on a detection threshold. We find that the first LQG-BH model produces significantly stronger signatures in EMRI signals than the second, making its quantum gravity effects more accessible to future space-borne gravitational-wave detection.
@@ I Introduction
Since its formulation, Einstein’s general relativity (GR) has provided an exceptionally successful framework for gravitation and the large-scale structure of spacetime Einstein:1915ca ; Einstein:1916vd . To date, it has withstood virtually all existing experimental and observational tests, from solar system dynamics Will:2014kxa ; GRAVITY:2020gka to the recent detections of gravitational waves (GWs) LIGOScientific:2016aoc ; LIGOScientific:2016lio ; LIGOScientific:2016sjg ; LIGOScientific:2019fpa ; LIGOScientific:2025rid and the images of Messier 87 (M 87) or Sagittarius A (Sgr A) EventHorizonTelescope:2019dse ; EventHorizonTelescope:2019ths ; EventHorizonTelescope:2022xnr ; EventHorizonTelescope:2022xqj , the latter two also confirmed the existence of black holes (BHs). Despite these successes, GR continues to encounter several fundamental theoretical challenges. A prominent example is the singularity theorems, formulated by Penrose and Hawking in the 1970s Penrose:1964wq ; Penrose:1969pc ; Hawking:1970zqf , which demonstrate the generic inevitable occurrence of spacetime singularities within GR. The existence of singularities leads to the non-extendibility of the spacetime manifold and to the geodesic incompleteness of timelike or null trajectories. In such extreme regimes, the failure of the classical description of gravity and the breakdown of known physical laws motivate the search for a more fundamental framework. This pursuit has led to the development of quantum gravity theories, such as loop quantum gravity (LQG) rovelli2004quantum ; Ashtekar:2004eh , which aim to provide a consistent description of spacetime and gravitation at the Planck scale, and have become an important focus of contemporary research in cosmology and BH physics.
@@
Given that spacetime covariance is a cornerstone of GR, its consistent implementation has been a primary theoretical pursuit in the development of LQG-BH models, with notable progress recently achieved (see Refs. Alonso-Bardaji:2021yls ; Alonso-Bardaji:2022ear ). In particular, building on the general scheme proposed in Bojowald:2011aa , C. Zhang et al. constructed the corresponding effective metric and Hamiltonian constraint while explicitly respecting covariance Zhang:2024khj . By applying two distinct polymerization schemes to the Dirac observable, i.e. the effective mass, they obtained two respective forms of the Hamiltonian constraint. Following the procedure outlined in Cafaro:2024vrw , the resulting dynamics were solved, yielding two classes of static, spherically symmetric BH solutions. In the first class, the metric components satisfy $g_{00}=g^{11}$, leading to a Reissner-Nordström (RN)-like structure. The second class features a metric whose temporal component coincides with that of the Schwarzschild solution, while the radial component $g_{11}$ contains a nontrivial correction function. In both cases, the quantum gravity effects are characterized by the parameter $\zeta$, which removes the classical singularity through a bounce region, resulting in a symmetric spacetime connecting to a white hole on the opposite side. These two families of LQG-BH solutions have subsequently been studied extensively across various physical contexts. For detailed investigations into their properties, perturbations, observational signatures, and related phenomenological implications, we refer the reader to Refs. Yang:2025ufs ; Konoplya:2024lch ; Skvortsova:2024msa ; Liu:2024soc ; Liu:2024wal ; Du:2024ujg ; Xamidov:2025oqx ; Umarov:2025wzm ; Chen:2025aqh ; Ban:2024qsa ; Lin:2024beb ; Heidari:2024bkm ; Wang:2024iwt ; Shu:2024tut ; Chen:2025ifv ; Du:2025kcx ; Liu:2025hcx ; Al-Badawi:2025yqu ; Zhang:2025ccx ; Sahlmann:2025fde ; Liu:2025iby .
@@
As indicated in previous studies Yang:2025ufs ; Konoplya:2024lch ; Skvortsova:2024msa ; Liu:2024soc ; Liu:2024wal ; Du:2024ujg ; Xamidov:2025oqx ; Umarov:2025wzm ; Chen:2025aqh ; Ban:2024qsa ; Lin:2024beb ; Heidari:2024bkm ; Wang:2024iwt ; Shu:2024tut ; Chen:2025ifv ; Du:2025kcx ; Liu:2025hcx ; Al-Badawi:2025yqu ; Zhang:2025ccx ; Sahlmann:2025fde ; Liu:2025iby , these quantum gravity effects may exist outside the horizon and could, in principle, be detectable through strong-field GW observations. Among various GW sources, extreme mass-ratio inspirals (EMRIs) stand out as high-precision laboratories for testing fundamental gravity, owing to their extreme mass ratios, high expected event rates, and large signal-to-noise ratios (SNRs) Amaro-Seoane:2007osp ; Cardenas-Avendano:2024mqp . A typical EMRI system consists of a stellar-mass compact secondary object orbiting and gradually inspiraling toward a supermassive black hole, emitting GWs at the same time, with mass ratios generally ranging from $10^{-7}$ to $10^{-4}$. By the late stages of inspiral, the secondary can complete on the order of $10^{4}$ to $10^{5}$ orbital cycles. This long-duration, cumulative evolution allows EMRI waveforms to encode exceptionally rich information about the strong-field spacetime, making them one of the most promising sources for future space-borne GW detectors such as LISA Babak:2017tow ; LISA:2022kgy , Taiji Gong:2021gvw , and TianQin TianQin:2020hid . At the same time, the intricate nature of these signals poses substantial challenges for accurate waveform modeling and data analysis. High fidelity EMRI waveform construction requires accounting for gravitational self-force effects Barack:2018yvs , which have so far been developed up to second order in Schwarzschild spacetime Wardell:2021fyy . Such calculations, however, are extremely complex and computationally expensive. Consequently, semi-classical kludge approximations — including the analytic kludge (AK) Barack:2003fp , numerical kludge (NK) Babak:2006uv , and augmented analytic kludge (AAK) Chua:2017ujo ; Katz:2021yft — are widely used as efficient waveform templates. These models reproduce the main physical features of realistic EMRI signals while greatly reducing computational cost.
@@
EMRIs have been applied to a wide range of research topics. They have been used, for instance, to test modified gravity theories, including those involving dark matter Dai:2023cft ; Zhang:2024ugv or scalar fields Guo:2022euk ; Barsanti:2022vvl ; Zhang:2022rfr , and to probe possible deviations from GR. They also serve as standard sirens for constraining the cosmological parameters Laghi:2021icm ; Laghi:2021pqk . In addition, they have inspired the development of specialized subfields, such as studies of dirty EMRIs, which focus on inspirals occurring in environments with accretion and matter around galactic-center black holes Lyu:2024gnk , and binary EMRIs ($b$-EMRIs), in which a binary system inspirals as a whole toward a supermassive black hole Chen:2018axp . More recently, EMRI systems have also been proposed as probes of quantum gravity effects Fu:2024cfk ; Liu:2024qci ; Yang:2024cnd ; Yang:2025esa ; Gong:2025mne . These works suggest that GW signals from EMRIs could carry subtle imprints of quantum gravity, and that EMRI observations can place tighter constraints on LQG parameters than those obtained from weak-field solar system experiments Liu:2022qiz ; Chen:2023bao ; Ai:2025myf , highlighting EMRIs as powerful tools for exploring the quantum aspects of gravity.
@@
Following this framework, we investigate quantum gravity effects on the external spacetime by employing EMRI system around two aforementioned types of covariant LQG black holes. The corresponding EMRI gravitational wave signals are generated using an AAK waveform model improved with the FastEMRIWaveforms (FEW) scheme Katz:2021yft . We then evaluate the detectability of such quantum gravity effects with LISA and derive constraints on the fundamental correction parameter $\zeta$. This study aims to quantify the deviations of LQG black holes from their GR counterparts and to compare the differences between the two LQG spacetimes, thereby providing potential waveform templates for future observations and contributing an essential step in theoretical testing.
@@
The paper is organized as follows. In Sec. II, we briefly introduce the spacetime properties of the two LQG black hole solutions and present the general equations of motion for a secondary body orbiting either primary BH in an EMRI setup. In Sec. III, we formulate the adiabatic evolution of the orbit under gravitational radiation, derive the associated energy and angular momentum fluxes, and present the resulting time evolution equations. The generation of EMRI gravitational waveforms and their detectability by the LISA mission are discussed in Sec. IV. Finally, Sec. V summarizes our findings and discusses future prospects.
@@
II Motions of massive particles in effective LQG black holes
In this section, we briefly review the spacetime geometrical properties of the two LQG black hole backgrounds derived from the condition of general covariance Zhang:2024khj . We then derive the universal equations of motion for a secondary compact object orbiting these central primary LQG black holes in the context of an EMRI, laying the groundwork for the subsequent inspiral evolution analysis.
II.1 Effective LQG black holes
Under the requirement of general covariance, the authors obtained an effective Hamiltonian constraint formalism and, via specialized polymerizations of Dirac observables, yielded two effective black hole metrics parametrized by a quantum correction. The spherically symmetric exteriors of these two LQG black holes can be described by the following general expression Zhang:2024khj :
@@ II.1 Effective LQG black holes
The metric functions for the two types are given separately. For the type-I black hole:
@@
Prior to advancing, we offer insights into the global geometric properties of these effective LQG black holes. For the type-I metric (2), the RN-like formalism gives two horizons where the outer one is the classical black hole horizon locates at $r_{+}=2M$, whereas the inner horizon $r_{-}$ is located at:
@@
For the type-II solution, the bounce surface coincides with the inner horizon, $r_{\mathrm{b}}=r_{-}$, serving as the non-singular bridge between the black hole and a white hole spacetime (see the supplementary material of Ref. Zhang:2024khj for more detailed discussions). We analyze the nature of this surface using the Kretschmann scalar, which diverges at a true singularity but remains finite at a coordinate singularity. Denoting the Kretschmann scalar of this LQG spacetime as $K_{\mathrm{LQG-II}}$, the computed expression is given below:
@@
Figure 1: Relation between $K_{\mathrm{LQG-II}}$ and the LQG parameter $\zeta$ at the bounce radius $r_{\mathrm{b}}$ in the second type LQG black hole background. The parameter $\zeta$ is varied within the range $\left[0,0.5\right]$.
@@ II.2 Geodesic for massive particle
In the EMRI system, we can approximate the secondary compact object as a test particle of mass $m$. Its dynamical evolution in the LQG black hole spacetime is derived from the following Lagrangian:
@@ III.1 Eccentric motion and fundamental orbital frequencies
For the marginally bound orbits (MBOs), the conditions $V_{\mathrm{eff}}=E^{2}/m^{2}$ and $\partial_{r}V_{\mathrm{eff}}|_{r_{\mathrm{p}}}=0$ must be satisfied. These lead to two real root solutions for each of the two types of LQG-BHs, which are given by:
@@
The first solution for $p$ (Eq. (32)) is constrained to the range $p\in\left(6+2e,\infty\right)$. This corresponds to a periastron radius $r_{\rm{p}}>{\left(6+2e\right)M}/\left(1+e\right)$, which implies $p>6$ and $r_{\rm{p}}>4M$ Cutler:1994pb ; Sundararajan:2008bw . The orbit becomes a plunging trajectory once the boundary $p=6+2e$ is crossed, making it the natural cutoff for our subsequent analysis. The second solution for $p$ (Eq. (33)) is beyond our study, as it is a consequence of the transition region involving quantum gravity and resides within the black hole, thus we do not consider it further here.
@@ III.2 Fluxes and Evolution
It should be noted that the exact expressions are intricate. Therefore, the expansions above are provided primarily to present our main results clearly. As shown, the dominant negative sign on the right-hand side indicates that the orbital parameters $\left\{p,e\right\}$ decrease with time — behavior consistent with the inspiral of the secondary body toward the central black hole. Quantum gravity corrections enter through higher-order terms and consequently modify the orbital evolution. The corresponding expressions $\left\{\dot{p},\dot{e}\right\}_{\mathrm{I}}$ for the LQG-BH I background can be derived following the same procedure.
@@
To illustrate the influence of the quantum parameter on the orbital dynamics within the two LQG-BH spacetimes, we numerically solve for the time evolution of $p\left(t\right)$ and $e\left(t\right)$ using the NDSolve routine in $\mathit{Mathematica}$, with initial conditions $\left\{{p}_{0},{e}_{0}\right\}$. According to the discussion in Sec. III.1, the separatrix of the orbital evolution in EMRI is located at $p=6+2e$. To maintain numerical stability, we terminate the integration when the orbit reaches $p_{\mathrm{stop}}=6+2e+0.1$, i.e., 0.1 away from the separatrix. The initial semi-latus rectum is fixed at $p_{0}=10$, providing at least one year of stable GW evolution within the interval $p\in\left[p_{\mathrm{stop}},\ p_{0}\right]$, and enabling a clear assessment of the parameter $\zeta$’s impacts on the waveform properties. For the initial eccentricity $e_{0}$, we consider values of 0.01 and 0.1. This choice is motivated by our strategy of ensuring that the orbital evolution remains sufficiently far from the circular bound orbit, corresponding to low-eccentricity elliptical orbits that retain moderate ellipticity without circularizing completely during the evolution. Under this setup, we concentrate on the gravitational-wave emission characteristics of low-eccentricity elliptical orbits.
@@
As the correction associated with the quantum parameter $\zeta$ arises only at higher orders, we separately compute the time evolution of the orbital parameter differences between the Schwarzschild spacetime (i.e., $\zeta=0$) and the two LQG-BH spacetimes for various choice of $\zeta$, the corresponding results are presented on a log-log scale in Figs. 2 and 3. For the calculations, we adopt masses of $m=10M_{\odot}$ for the secondary body and $M=10^{6}M_{\odot}$ for the central object, with $M_{\odot}$ denoting the solar mass. The numerical results show that the orbital deviations $\Delta p$ and $\Delta e$ for both types of LQG black holes exhibit a monotonic increase over time, with their magnitudes growing as $\zeta$ increases. This behavior indicates that quantum gravity effects induce a cumulative impact on orbital evolution, thereby offering potential observability over long timescales. Moreover, a comparison between the two LQG-BH models reveals that for a given $\zeta$, the orbital parameter deviations in the LQG-BH I spacetime are more pronounced than those in LQG-BH II. However, the evolution is only weakly affected by the choice of initial conditions. In other words, the orbital parameter differences in LQG-BH I are more sensitive to variations in $\zeta$ but less sensitive to the initial configuration.
@@
Figure 2: Deviation in the time evolution of the semi-latus rectum $\Delta p=p_{\mathrm{I,II}}-p_{\mathrm{Sch}}$ for different values of the quantum gravity parameter $\zeta$. The first row corresponds to the LQG-BH I background, and the second row to the LQG-BH II background. The left column shows results for the initial eccentricity $e_{0}=0.01$, and the right column for $e_{0}=0.1$. The initial semi-latus rectum is set to $p_{0}=10$ in all cases.
@@
Figure 3: Deviation in the time evolution of the eccentricity $\Delta e=e_{\mathrm{I,II}}-e_{\mathrm{Sch}}$ for different values of the quantum gravity parameter $\zeta$. The first row corresponds to the LQG-BH I background, and the second row to the LQG-BH II background. The left column shows results for the initial eccentricity $e_{0}=0.01$, and the right column for $e_{0}=0.1$. The initial semi-latus rectum is set to $p_{0}=10$ in all cases.
@@
Evidently, the phase $\Phi_{\phi,r}\left(t\right)$ depends entirely on the orbital evolution parameters $\left\{p\left(t\right),e\left(t\right)\right\}$. Therefore, we still analyze the effects of the parameter $\zeta$ by numerically computing the relative phase deviation between cases with and without quantum gravity effects. Fig. 4 shows the evolution of the dephasing $\left|\Delta\Phi\right|\equiv\left|{\Phi_{\phi}}_{\mathrm{I,II}}-{\Phi_{\phi}}_{\mathrm{Sch}}\right|$ over time for different $\zeta$ values in these two black hole models, with both initial phase set to 0. The figure clearly shows that the phase difference grows with time for all chosen values of $\zeta$, and the departure from the Schwarzschild case becomes increasingly evident as $\zeta$ rises. This behavior indicates that the accumulated orbital phases may carry discernible imprints of quantum gravity effects in the GW signal. At fixed $\zeta$, the discrepancy is considerably stronger in the LQG-BH I spacetime, whereas the variation in the LQG-BH II background remains comparatively weak. These results suggest that LQG-BH I offers more favorable conditions for revealing quantum gravity signatures in EMRI waveforms, while the LQG-BH II scenario would require larger $\zeta$ to yield similarly measurable deviations.
Figure 4: EMRI orbital dephasing $\left|\Delta\Phi\right|=\left|{\Phi_{\phi}}_{\mathrm{I,II}}-{\Phi_{\phi}}_{\mathrm{Sch}}\right|$ versus evolution time for different values of $\zeta$, where the inset presents enlarged views for $\zeta=1/1000$ and $\zeta=1/100$. The first row corresponds to the LQG-BH I background, and the second row to the LQG-BH II background. The left column shows results for the initial eccentricity $e_{0}=0.01$, and the right column for $e_{0}=0.1$. The fixed initial parameters $\left\{M,m,p_{0},{\Phi_{\phi}}_{0},{\Phi_{r}}_{0}\right\}$ are set to $\left\{10^{6}M_{\odot},10M_{\odot},10,0,0\right\}$ correspondingly.

### v2

@@ Abstract
Within a generally covariant Hamiltonian framework of loop quantum gravity (LQG), two black hole models parameterized by a quantum correction $\zeta$ have recently been constructed. Using extreme mass-ratio inspirals (EMRIs) as high-precision probes, we investigate the imprints of this LQG deformation in the surrounding spacetime. Waveforms generated via an improved augmented analytic kludge (AAK) model in both LQG black hole backgrounds and in Schwarzschild spacetime are compared through a faithfulness analysis. This allows us to quantify the detectability of the deviation with LISA and to derive constraints on $\zeta$ based on a detection threshold. We find that the first LQG black hole model produces significantly stronger signatures in EMRI signals than the second, making its quantum gravity effects more accessible to future space-borne gravitational-wave detection.
@@ I Introduction
Since its formulation, Einstein’s general relativity (GR) has provided an exceptionally successful framework for gravitation and the large-scale structure of spacetime Einstein:1915ca ; Einstein:1916vd . To date, it has withstood virtually all existing experimental and observational tests, from solar system dynamics Will:2014kxa ; GRAVITY:2020gka to the recent detections of gravitational waves (GWs) LIGOScientific:2016aoc ; LIGOScientific:2016lio ; LIGOScientific:2016sjg ; LIGOScientific:2019fpa ; LIGOScientific:2025rid and the images of M 87 or Sgr A EventHorizonTelescope:2019dse ; EventHorizonTelescope:2019ths ; EventHorizonTelescope:2022xnr ; EventHorizonTelescope:2022xqj , the latter two also confirmed the existence of black holes (BHs). Despite these successes, GR continues to encounter several fundamental theoretical challenges. A prominent example is the singularity theorems, formulated by Penrose and Hawking in the 1970s Penrose:1964wq ; Penrose:1969pc ; Hawking:1970zqf , which demonstrate the generic inevitable occurrence of spacetime singularities within GR. The existence of singularities leads to the non-extendibility of the spacetime manifold and to the geodesic incompleteness of timelike or null trajectories. In such extreme regimes, the failure of the classical description of gravity and the breakdown of known physical laws motivate the search for a more fundamental framework. This pursuit has led to the development of quantum gravity theories, such as loop quantum gravity (LQG) rovelli2004quantum ; Ashtekar:2004eh , which aim to provide a consistent description of spacetime and gravitation at the Planck scale, and have become an important focus of contemporary research in cosmology and BH physics.
@@
Given that spacetime covariance is a cornerstone of GR, its consistent implementation has been a primary theoretical pursuit in the development of the loop quantum gravity black hole (LQG-BH) models, with notable progress recently achieved (see Refs. Alonso-Bardaji:2021yls ; Alonso-Bardaji:2022ear ). In particular, building on the general scheme proposed in Bojowald:2011aa , C. Zhang et al. constructed the corresponding effective metric and Hamiltonian constraint while explicitly respecting covariance Zhang:2024khj . By applying two distinct polymerization schemes to the Dirac observable, i.e. the effective mass, they obtained two respective forms of the Hamiltonian constraint. Following the procedure outlined in Cafaro:2024vrw , the resulting dynamics were solved, yielding two classes of static, spherically symmetric BH solutions. In the first class, the metric components satisfy $g_{00}=g^{11}$, leading to a Reissner-Nordström (RN)-like structure. The second class features a metric whose temporal component coincides with that of the Schwarzschild solution, while the radial component $g_{11}$ contains a nontrivial correction function. In both cases, the quantum gravity effects are characterized by the parameter $\zeta$, which removes the classical singularity through a bounce region, resulting in a symmetric spacetime connecting to a white hole on the opposite side. These two families of LQG-BH solutions have subsequently been studied extensively across various physical contexts. For detailed investigations into their properties, perturbations, observational signatures, and related phenomenological implications, we refer the reader to Refs. Yang:2025ufs ; Konoplya:2024lch ; Skvortsova:2024msa ; Liu:2024soc ; Liu:2024wal ; Du:2024ujg ; Xamidov:2025oqx ; Umarov:2025wzm ; Chen:2025aqh ; Ban:2024qsa ; Lin:2024beb ; Heidari:2024bkm ; Wang:2024iwt ; Shu:2024tut ; Chen:2025ifv ; Du:2025kcx ; Liu:2025hcx ; Al-Badawi:2025yqu ; Zhang:2025ccx ; Sahlmann:2025fde ; Liu:2025iby .
@@
As indicated in previous studies Yang:2025ufs ; Konoplya:2024lch ; Skvortsova:2024msa ; Liu:2024soc ; Liu:2024wal ; Du:2024ujg ; Xamidov:2025oqx ; Umarov:2025wzm ; Chen:2025aqh ; Ban:2024qsa ; Lin:2024beb ; Heidari:2024bkm ; Wang:2024iwt ; Shu:2024tut ; Chen:2025ifv ; Du:2025kcx ; Liu:2025hcx ; Al-Badawi:2025yqu ; Zhang:2025ccx ; Sahlmann:2025fde ; Liu:2025iby , these quantum gravity effects may exist outside the horizon and could, in principle, be detectable through strong-field GW observations. GWs provide one of the most precise probes to study the detailed structure of BH spacetimes and their surrounding astrophysical environment Nair:2022xfm ; Mitra:2023sny ; Krishnendu:2025byo ; Battista:2021rlh ; Battista:2022hmv ; DeFalco:2023djo ; DeFalco:2024ojf . Among various GW sources, extreme mass-ratio inspirals (EMRIs) stand out as high-precision laboratories for testing fundamental gravity, owing to their extreme mass ratios, high expected event rates, and large signal-to-noise ratios (SNRs) Amaro-Seoane:2007osp ; Cardenas-Avendano:2024mqp . A typical EMRI system consists of a stellar-mass compact secondary object orbiting and gradually inspiraling toward a supermassive BH, emitting GWs at the same time, with mass ratios generally ranging from $10^{-7}$ to $10^{-4}$. By the late stages of inspiral, the secondary can complete on the order of $10^{4}$ to $10^{5}$ orbital cycles. This long-duration, cumulative evolution allows EMRI waveforms to encode exceptionally rich information about the strong-field spacetime, making them one of the most promising sources for future space-borne GW detectors such as LISA Babak:2017tow ; LISA:2022kgy , Taiji Gong:2021gvw , and TianQin TianQin:2020hid . At the same time, the intricate nature of these signals poses substantial challenges for accurate waveform modeling and data analysis. High fidelity EMRI waveform construction requires accounting for gravitational self-force effects Barack:2018yvs , which have so far been developed up to second order in Schwarzschild spacetime Wardell:2021fyy . Such calculations, however, are extremely complex and computationally expensive. Consequently, semi-classical kludge approximations — including the analytic kludge (AK) Barack:2003fp , numerical kludge (NK) Babak:2006uv , and augmented analytic kludge (AAK) Chua:2017ujo ; Katz:2021yft — are widely used as efficient waveform templates. These models reproduce the main physical features of realistic EMRI signals while greatly reducing computational cost.
@@
EMRIs have been applied to a wide range of research topics. They have been used, for instance, to test modified gravity theories, including those involving dark matter Dai:2023cft ; Zhang:2024ugv or scalar fields Guo:2022euk ; Barsanti:2022vvl ; Zhang:2022rfr ; Zi:2025lio , and to probe possible deviations from GR Kumar:2024utz ; Ahmed:2025azu . They also serve as standard sirens for constraining the cosmological parameters Laghi:2021icm ; Laghi:2021pqk . In addition, they have inspired the development of specialized subfields, such as studies of dirty EMRIs, which focus on inspirals occurring in environments with accretion and matter around galactic-center BHs Lyu:2024gnk , and binary EMRIs, in which a binary system inspirals as a whole toward a supermassive BH Chen:2018axp . More recently, EMRI systems have also been proposed as probes of quantum gravity effects Fu:2024cfk ; Liu:2024qci ; Zi:2024jla ; Yang:2024cnd ; Yang:2025esa ; Gong:2025mne ; Ahmed:2025shr . These works suggest that GW signals from EMRIs could carry subtle imprints of quantum gravity, and that EMRI observations can place tighter constraints on LQG parameters than those obtained from weak-field solar system experiments Liu:2022qiz ; Chen:2023bao ; Ai:2025myf , highlighting EMRIs as powerful tools for exploring the quantum aspects of gravity.
@@
Following this framework, we investigate quantum gravity effects on the external spacetime by employing EMRI system around two aforementioned types of covariant LQG-BHs. The corresponding EMRI signals are generated using an AAK waveform model improved with the FastEMRIWaveforms (FEW) scheme Katz:2021yft . We then evaluate the detectability of such quantum gravity effects with LISA and derive constraints on the fundamental correction parameter $\zeta$. This study aims to quantify the deviations of LQG-BHs from their GR counterparts and to compare the differences between the two LQG spacetimes, thereby providing potential waveform templates for future observations and contributing an essential step in theoretical testing.
@@
The paper is organized as follows. In Sec. II, we briefly introduce the spacetime properties of the two LQG-corrected BH solutions and present the general equations of motion for a secondary body orbiting either primary BH in an EMRI setup. In Sec. III, we formulate the adiabatic evolution of the orbit under gravitational radiation, derive the associated energy and angular momentum fluxes, and present the resulting time evolution equations. The generation of EMRI gravitational waveforms and their detectability by the LISA mission are discussed in Sec. IV. Finally, Sec. V summarizes our findings and discusses future prospects.
@@
II Motions of massive particles in effective LQG-BHs
In this section, we briefly review the spacetime geometrical properties of the two LQG-BH backgrounds derived from the condition of general covariance Zhang:2024khj . We then derive the universal equations of motion for a secondary compact object orbiting these central primary LQG-BHs in the context of an EMRI, laying the groundwork for the subsequent inspiral evolution analysis.
II.1 Effective LQG-BHs
Under the requirement of general covariance, the authors obtained an effective Hamiltonian constraint formalism and, via specialized polymerizations of Dirac observables, yielded two effective BH metrics parametrized by a quantum correction. The spherically symmetric exteriors of these two LQG-BHs can be described by the following general expression Zhang:2024khj :
@@ II.1 Effective LQG-BHs
The metric functions for the two types are given separately. For the type-I BH:
@@
Prior to advancing, we offer insights into the global geometric properties of these effective LQG-based BHs. For the type-I metric (2), the RN-like formalism gives two horizons where the outer one is the classical BH horizon locates at $r_{+}=2M$, whereas the inner horizon $r_{-}$ is located at:
@@
For the type-II solution, the bounce surface coincides with the inner horizon, $r_{\mathrm{b}}=r_{-}$, serving as the non-singular bridge between the BH and a white hole spacetime (see the supplementary material of Ref. Zhang:2024khj for more detailed discussions). We analyze the nature of this surface using the Kretschmann scalar, which diverges at a true singularity but remains finite at a coordinate singularity. Denoting the Kretschmann scalar of this LQG spacetime as $K_{\mathrm{LQG-II}}$, the computed expression is given below:
@@
Figure 1: Relation between $K_{\mathrm{LQG-II}}$ and the LQG parameter $\zeta$ at the bounce radius $r_{\mathrm{b}}$.
@@ II.2 Geodesic for massive particle
In the EMRI system, we can approximate the secondary compact object as a test particle of mass $m$. Its dynamical evolution in the LQG-BH spacetime is derived from the following Lagrangian:
@@ III.1 Eccentric motion and fundamental orbital frequencies
For the marginally bound orbits, the conditions $V_{\mathrm{eff}}=E^{2}/m^{2}$ and $\partial_{r}V_{\mathrm{eff}}|_{r_{\mathrm{p}}}=0$ must be satisfied. These lead to two real root solutions for each of the two types of LQG-BHs, which are given by:
@@
The first solution for $p$ (Eq. (32)) is constrained to the range $p\in\left(6+2e,\infty\right)$. This corresponds to a periastron radius $r_{\rm{p}}>{\left(6+2e\right)M}/\left(1+e\right)$, which implies $p>6$ and $r_{\rm{p}}>4M$ Cutler:1994pb ; Sundararajan:2008bw . The orbit becomes a plunging trajectory once the boundary $p=6+2e$ is crossed, making it the natural cutoff for our subsequent analysis. The second solution for $p$ (Eq. (33)) is beyond our study, as it is a consequence of the transition region involving quantum gravity and resides within the BH, thus we do not consider it further here.
@@ III.2 Fluxes and evolution
It should be noted that the exact expressions are intricate. Therefore, the expansions above are provided primarily to present our main results clearly. As shown, the dominant negative sign on the right-hand side indicates that the orbital parameters $\left\{p,e\right\}$ decrease with time — behavior consistent with the inspiral of the secondary body toward the central BH. Quantum gravity corrections enter through higher-order terms and consequently modify the orbital evolution. The corresponding expressions $\left\{\dot{p},\dot{e}\right\}_{\mathrm{I}}$ for the LQG-BH I background can be derived following the same procedure.
@@
To illustrate the influence of the quantum parameter on the orbital dynamics within the two LQG-BH spacetimes, we numerically solve for the time evolution of $p\left(t\right)$ and $e\left(t\right)$ using the NDSolve routine in $\mathit{Mathematica}$, with initial conditions $\left\{{p}_{0},{e}_{0}\right\}$. According to the discussion in subsection III.1, the separatrix of the orbital evolution in EMRI is located at $p=6+2e$. To maintain numerical stability, we terminate the integration when the orbit reaches $p_{\mathrm{stop}}=6+2e+0.1$, i.e., 0.1 away from the separatrix. The initial semi-latus rectum is fixed at $p_{0}=10$, providing at least one year of stable GW evolution within the interval $p\in\left[p_{\mathrm{stop}},\ p_{0}\right]$, and enabling a clear assessment of the parameter $\zeta$’s impacts on the waveform properties. For the initial eccentricity $e_{0}$, we consider values of 0.01 and 0.1. This choice is motivated by our strategy of ensuring that the orbital evolution remains sufficiently far from the circular bound orbit, corresponding to low-eccentricity elliptical orbits that retain moderate ellipticity without circularizing completely during the evolution. Under this setup, we concentrate on the GW emission characteristics of low-eccentricity elliptical orbits.
@@
As the correction associated with the quantum parameter $\zeta$ arises only at higher orders, we separately compute the time evolution of the orbital parameter differences between the Schwarzschild spacetime (i.e., $\zeta=0$) and the two LQG-BH spacetimes for various choice of $\zeta$, the corresponding results are presented on a log-log scale in Figs. 2 and 3. For the calculations, we adopt masses of $m=10M_{\odot}$ for the secondary body and $M=10^{6}M_{\odot}$ for the central object, with $M_{\odot}$ denoting the solar mass. The numerical results show that the orbital deviations $\Delta p$ and $\Delta e$ for both types of LQG-BHs exhibit a monotonic increase over time, with their magnitudes growing as $\zeta$ increases. This behavior indicates that quantum gravity effects induce a cumulative impact on orbital evolution, thereby offering potential observability over long timescales. Moreover, a comparison between the two LQG-BH models reveals that for a given $\zeta$, the orbital parameter deviations in the LQG-BH I spacetime are more pronounced than those in LQG-BH II. However, the evolution is only weakly affected by the choice of initial conditions. In other words, the orbital parameter differences in LQG-BH I are more sensitive to variations in $\zeta$ but less sensitive to the initial configuration.
@@
Figure 2: Deviation in the time evolution of the semi-latus rectum $\Delta p=p_{\mathrm{I,II}}-p_{\mathrm{Sch}}$ for different values of the quantum gravity parameter $\zeta$.
@@
Figure 3: Deviation in the time evolution of the eccentricity $\Delta e=e_{\mathrm{I,II}}-e_{\mathrm{Sch}}$ for different values of the quantum gravity parameter $\zeta$.
@@
Evidently, the phase $\Phi_{\phi,r}\left(t\right)$ depends entirely on the orbital evolution parameters $\left\{p\left(t\right),e\left(t\right)\right\}$. Therefore, we still analyze the effects of the parameter $\zeta$ by numerically computing the relative phase deviation between cases with and without quantum gravity effects. Fig. 4 shows the evolution of the dephasing $\left|\Delta\Phi\right|\equiv\left|{\Phi_{\phi}}_{\mathrm{I,II}}-{\Phi_{\phi}}_{\mathrm{Sch}}\right|$ over time for different $\zeta$ values in these two BH models, with both initial phase set to 0. The figure clearly shows that the phase difference grows with time for all chosen values of $\zeta$, and the departure from the Schwarzschild case becomes increasingly evident as $\zeta$ rises. This behavior indicates that the accumulated orbital phases may carry discernible imprints of quantum gravity effects in the GW signal. At fixed $\zeta$, the discrepancy is considerably stronger in the LQG-BH I spacetime, whereas the variation in the LQG-BH II background remains comparatively weak. These results suggest that LQG-BH I offers more favorable conditions for revealing quantum gravity signatures in EMRI waveforms, while the LQG-BH II scenario would require larger $\zeta$ to yield similarly measurable deviations.
Figure 4: EMRI orbital dephasing $\left|\Delta\Phi\right|$ versus evolution time for different values of $\zeta$, where the inset presents enlarged views for $\zeta=1/1000$ and $\zeta=1/100$.